A 5.5 g bullet moving at 120 m/s strikes a thick wall. Assume that the bullet


undergoes a uniform deceleration and stops after penetrating 6 cm within the wall. Find


(a) The time taken by the bullet to stop.


(b) The impulse on the wall.


(c) The magnitude of the average force experienced by the wall

1
Expert's answer
2021-12-16T11:33:14-0500

(a) "S=v^2/(2a)\to a=v^2/(2S)=120^2/(2\cdot0.06)="


"=120000\ (m/s^2)"


"v=v_0-at\to t=v_0/a=120/120000=0.001\ (s)"


(b)"\ \ p=mv=0.0055\cdot120=0.66\ (kg\cdot m/s)"


(c) "F=\Delta p/\Delta t=0.66/0.001=660\ (N)"